OpenStudy (anonymous):

How to find the domain and range of the function A(x)=4x^2+1 without graphing?

OpenStudy (jdoe0001):

well the domain is simple, \(\bf \mathbb{R}\) now the range keep in mind that, for any "x" you throw in, it'll be squared, that means including any negative values for "x", -25, -1000 and so on once squared, it will never give you a negative number, so we can say "y" will never be negative the lowest values \(\bf x^2\) can give you is 0, and when that happens \(\bf 4(0)^2+1 \implies 1\) that's the lowest "y" will get, and then it goes right back up

OpenStudy (anonymous):

so, is the domain \[(1,\infty)\]

OpenStudy (jdoe0001):

well, the domain will be \(\bf \mathbb{R} \implies (-\infty, +\infty)\)

OpenStudy (anonymous):

okay, will the range be (1, infinty)

OpenStudy (jdoe0001):

yes, the domain is not constrained by anything, the range is

OpenStudy (anonymous):

okay that makes sense. thankyou so much.

OpenStudy (jdoe0001):

yw